OpenStudy (anonymous):

Simplify the following expression. √-14√-56

OpenStudy (anonymous):

-28

OpenStudy (anonymous):

- 28

OpenStudy (angela210793):

isqr14*isqr56=i^2sqr(14*56)=-sqr(784)=-28